OpenStudy (anonymous):

Multiply: (6-5i)(4+3i)

OpenStudy (anonymous):

39-2i

OpenStudy (anonymous):

39-2i also

OpenStudy (anonymous):

24+18i-20i-15i^2=24-2i+15=39-2i

OpenStudy (anonymous):

Thank you!